# Extended Circular Nim

### v1

@@ 1.3.2 Selective sum
$\{G^{\prime}\vee H,G\vee H^{\prime},G^{\prime}\vee H^{\prime}\mid G^{\prime}\text{ is an option of }G,H^{\prime}\text{ is an option of }H\}.$
@@
2 Main results
2.1 Extended circular nim
Extended circular nim $ECN(m_{S},k)$ is defined as follows.
@@ 2.1 Extended circular nim
Let $m$ and $k$ be positive integers with $k\leq m$ and let $S$ be a set of positive integers that are not more than half of $m$. Extended circular nim ${\rm ECN}(m_{S},k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta=\bigcup_{s\in S}\bigcup_{i=0}^{m-1}\bigcup_{j=0}^{k-1}\{\{v_{i},v_{((i+s)\bmod m)},\ldots,v_{((i+js)\bmod m)}\}\}$.
That is, in extended circular nim $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$.
Figure 1 shows piles of a position in ${\rm ECN}(6_{\{1,2\}},3)$. Selecting piles $A,B,$ and $C$ is legal because the player selects every first pile (connected black lines), also, selecting piles $A,C$, and $E$ is legal because the player selects every second pile (connected green lines). However selecting piles $A,B$ , and $D$ is illegal in this ruleset.
Figure 1: ${\rm ECN}(6_{1,2},3)$
@@
Consider ${\rm ECN}(a_{S},a-1)$. If $b\in S$ is prime to $a$, the player can choose arbitrary $a-1$ piles. Therefore, ${\rm ECN}(a_{S},a-1)$ is isomorphic to ${\rm MN}(a,a-1)$. Otherwise, the player can choose at most $\frac{a}{2}$ piles. Since $\frac{a}{2}\leq a-2$, this ruleset is isomorphic to ${\rm ECN}(a_{S},a-2)$.
Similarly, ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm MN}(a,a)$ when an integer $b\in S$ is pritme to $a$ and otherwise ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm ECN}(a_{S},a-2)$. ∎
Table 1 summarizes these discussions. From them, when considering extended circular nim, we do not need to consider ${\rm ECN}(a_{S},1),{\rm ECN}(a_{S},a-1),{\rm ECN}(a_{S},a)$ independently. In this paper, we consider ${\rm ECN}(a_{S},i)(1<i<a-1)$.
@@ 2.2 Extended circular nim with less than six piles
Consider ${\rm ECN}(4_{S},2)$.
When $S=\{1\}$, the ruleset is already considered as circular nim.
When $S=\{2\},$ we can combine each diagonal piles as single piles. Therefore, the position can be considered as a position in two-pile nim.
Consier ${\rm ECN}(5_{S},k)(k=2,3)$. When $S=\{1\}$, the ruleset is already considered as circular nim. As shown in Fig. 2, ${\rm ECN}(5_{\{2\}},i)$ is isomorphic to ${\rm ECN}(5_{\{1\}},i)$. When $S=\{1,2\}$, arbitrary two piles can be selected in ${\rm ECN}(5_{\{1,2\}},2)$ and arbitrary three piles can be selected in ${\rm ECN}(5_{\{1,2\}},3)$. Therefore, these rulesets are isomorphic to ${\rm MN}(5,2)$ and ${\rm MN}(5,3)$, respectively.
Table 2 summarizes these resules.
@@
${\rm ECN}(4_{1},2)$ Shown in [2]
${\rm ECN}(4_{2},2)$ Regarded as $2$-pile nim
${\rm ECN}(4_{1,2},2)$ Isomorphic to ${\rm MN}(4,2)$
${\rm ECN}(5_{1},2)$ Shown in [4]
${\rm ECN}(5_{1},3)$ Shown in [4]
${\rm ECN}(5_{2},i)(2\leq i\leq 3)$ Isomorphic to ${\rm ECN}(5_{1},i)$
${\rm ECN}(5_{1,2},i)(2\leq i\leq 3)$ Isomorphic to ${\rm MN}(5,i)$
@@
2.3 Extended circular nim with six piles
@@
First, consider ${\rm ECN}(6_{\{2\}},2)$. This ruleset can be considered as a disjunctive sum of two positions in ${\rm MN}(3,2)$. In general, closed formula for Sprague-Grundy values of Moore’s nim is not found, but for the case of ${\rm MN}(n,n-1)$, the formula for Sprague-Grundy values is found in [7]. Thus, in ${\rm ECN}(6_{\{2\}},2)$, the set of $\mathcal{P}$-positions can be characterized.
@@
Let $N$ be the set of positions which are not in $P$. The terminal position is in $P$.
@@
Lastly, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in N$. Let $k=\min(n_{0}\oplus n_{3},n_{1}\oplus n_{4},n_{2}\oplus n_{5})$. Without loss of generality, $k=n_{0}\oplus n_{3}$. Then, at least one of $n_{1}\oplus n_{4}>k$ and $n_{2}\oplus n_{5}>k$ holds. In addition, each of $(n_{1}),(n_{4}),(n_{2}),(n_{5}),(n_{1},n_{2}),(n_{1},n_{5}),(n_{4},n_{2})$ and $(n_{4},n_{5})$ can be reduced in one move in ${\rm ECN}(6_{\{1,2\}},2)$. Therefore, there is an option $M^{\prime}$ of $M$ which satisfies $M^{\prime}=(n_{0},n^{\prime}_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n^{\prime}_{5})$ and $n_{0}\oplus n_{3}=n^{\prime}_{1}\oplus n^{\prime}_{4}=n^{\prime}_{2}\oplus n^{\prime}_{5}$. ∎
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}\}$.
@@
Consider a position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. This position is a selective sum of $(n_{0},n_{3})$, $(n_{1},n_{4})$, and $(n_{2},n_{5})$, which are positions of two-pile nim. Therefore, $n_{0}\oplus n_{3}=n_{1}\oplus n_{4}=n_{2}\oplus n_{5}=0$, that is, $n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}$ is a necessary and sufficient condition for $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position.
@@
Let $P$ be the set of positions which satisfies the following:
  • Assume that $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. When each $n_{i}$ is expanded in the binary form, the summands of each digit contain four or no $1$s. In addition, for each digit contain four $1$s, the $1$s are in $(n_{0},n_{1},n_{3},n_{4})$ or $(n_{0},n_{1},n_{2},n_{4})$.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},4)$. Then, M is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P$.
@@
This can be proven by using an early result on simplicial complex. We can use Theroem 4.4 in [4].
@@
Proof.
Consider a position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,3\}},2)$. This position is a selective sum of $(n_{0},n_{2},n_{4})$ and $(n_{1},n_{3},n_{5})$, which are positions of three-pile nim. Therefore, $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position if and only if $n_{0}\oplus n_{2}\oplus n_{4}=n_{1}\oplus n_{3}\oplus n_{5}=0$ ∎
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in $ECN(6_{\{2,3\}},3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{2}+n_{4}=n_{1}+n_{3}+n_{5},n_{0}\oplus n_{1}\oplus n_{2}=0,n_{0}\leq n_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}\}$.
@@
Next, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. Then, $n_{0}\leq n_{3},n_{1}\leq n_{4},$ and $n_{2}\leq n_{5}$ hold. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5})$ be an option of $M$. When the player reduces some of $n_{0},n_{2},n_{4}$, or reduces some of $n_{1},n_{3},n_{5}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}\neq n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ holds and $M^{\prime}\not\in P$. For the case that the player reduces the same number from $n_{0}$ and $n_{3}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ is satisfied but $n^{\prime}_{0}\oplus n^{\prime}_{1}\oplus n^{\prime}_{2}=n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$ holds. We show $M^{\prime}\not\in_{\circlearrowleft}P$ by contradiction:
@@
    Assume that $n^{\prime}_{2}\oplus n^{\prime}_{3}\oplus n^{\prime}_{4}=n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=0$. In order to satisfy $M^{\prime}\in P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4}$ hold. However, then we have $n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=n_{2}\oplus n^{\prime}_{0}\oplus n_{1}=0$, which contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.
@@
    Assume that $n^{\prime}_{3}\oplus n^{\prime}_{4}\oplus n^{\prime}_{5}=n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0$. In order to satisfy $M^{\prime}\in P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1},n_{5}\leq n_{2}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4},n_{2}=n_{5}$ hold. This contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.

### v2

@@ 1.3.2 Selective sum
[…]
@@
The rest of this paper is organized as follows: In Section 2, we introduce the rule of extended circular nim and show basic results. In Sections 3, 4, and 5, we present results on extended circular nim with six, seven, and eight piles, respectively. In these sections, we also prove Theorems 11, 14, and 16, which can be used in general cases.
2 Extended circular nim
Extended circular nim ${\rm ECN}(m_{S},k)$ is defined as follows.
@@ 2 Extended circular nim
Let $m$ and $k$ be positive integers with $k\leq m$ and let $S$ be a set of positive integers that are not more than half of $m$. Extended circular nim ${\rm ECN}(m_{S},k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta$ is the set of all non-empty subset of $\bigcup_{s\in S}\bigcup_{i=0}^{m-1}\{v_{i},v_{((i+s)\bmod m)},\ldots,v_{((i+(k-1)s)\bmod m)}\}$.
That is, in extended circular nim ${\rm ECN}(m_{S},k)$, there are $m$ piles of tokens arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$.
Figure 4 shows piles of a position in ${\rm ECN}(6_{\{1,2\}},3)$. For example, selecting piles $B,C,$ and $D$ is legal because the player selects every first pile (connected black lines), also, selecting piles $A,C$, and $E$ is legal because the player selects every second pile (connected green lines). However, selecting piles $A,B$, and $D$ is illegal in this ruleset.
Figure 4: ${\rm ECN}(6_{\{1,2\}},3)$.
@@
Consider ${\rm ECN}(a_{S},a-1)$. If $b\in S$ is prime to $a$, the player can choose arbitrary a1a-1 piles. Therefore, ${\rm ECN}(a_{S},a-1)$ is isomorphic to ${\rm MN}(a,a-1)$. Otherwise, the player can choose at most $\frac{a}{2}$ piles. Since $\frac{a}{2}\leq a-2$, this ruleset is isomorphic to ${\rm ECN}(a_{S},a-2)$.
Similarly, ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm MN}(a,a)$ when an integer $b\in S$ is prime to $a$ and otherwise ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm ECN}(a_{S},a-2)$. ∎
Table 1 summarizes these discussions. From them, when considering extended circular nim, we do not need to consider ${\rm ECN}(a_{S},1),{\rm ECN}(a_{S},a-1),{\rm ECN}(a_{S},a)$ independently. Thus, in this paper, we consider ${\rm ECN}(a_{S},i)(1<i<a-1)$.
@@ 2.1 Extended circular nim with less than six piles
First, we consider ${\rm ECN}(4_{S},2)$. When $S=\{1\}$, the ruleset is already considered as a case of circular nim in [6]. When $S=\{2\},$ we can regard each pair of diagonal piles as a single pile. Therefore, the position can be considered as a position in two-pile nim. When $S=\{1,2\},$ arbitrary two piles can be selected in ${\rm ECN}(4_{\{1,2\}},2)$. Therefore, this ruleset is isomorphic to ${\rm MN}(4,2).$
Next, we consider ${\rm ECN}(5_{S},k)(k=2,3)$. When $S=\{1\}$, the ruleset is already considered as cases of circular nim in [8, 5]. As shown in Fig. 5, ${\rm ECN}(5_{\{2\}},i)$ is isomorphic to ${\rm ECN}(5_{\{1\}},i)$. When $S=\{1,2\}$, arbitrary two piles can be selected in ${\rm ECN}(5_{\{1,2\}},2)$ and arbitrary three piles can be selected in ${\rm ECN}(5_{\{1,2\}},3)$. Therefore, these rulesets are isomorphic to ${\rm MN}(5,2)$ and ${\rm MN}(5,3)$, respectively.
Table 2 summarizes these results.
@@
${\rm ECN}(4_{\{1\}},2)$ Shown in [6]
${\rm ECN}(4_{\{2\}},2)$ Regarded as two-pile nim
${\rm ECN}(4_{\{1,2\}},2)$ Isomorphic to ${\rm MN}(4,2)$
${\rm ECN}(5_{\{1\}},i)(2\leq i\leq 3)$ Shown in [8, 5]
${\rm ECN}(5_{\{2\}},i)(2\leq i\leq 3)$ Isomorphic to ${\rm ECN}(5_{\{1\}},i)$
${\rm ECN}(5_{\{1,2\}},i)(2\leq i\leq 3)$ Isomorphic to ${\rm MN}(5,i)$
@@
3 Extended circular nim with six piles
@@
First, consider ${\rm ECN}(6_{\{2\}},2)$. This ruleset can be considered as a disjunctive sum of two positions in ${\rm MN}(3,2)$. Therefore, by using Theorem 5, the set of $\mathcal{P}$-positions can be characterized.
In ${\rm ECN}(6_{\{2\}},3)$ and ${\rm ECN}(6_{\{2\}},4)$, The position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ can be considered as a position $(n_{0}+n_{2}+n_{4},n_{1}+n_{3}+n_{5})$ in two-pile nim. Therefore, $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position if and only if $n_{0}+n_{2}+n_{4}=n_{1}+n_{3}+n_{5}.$ Also, in ${\rm ECN}(6_{\{3\}},2),{\rm ECN}(6_{\{3\}},3),$ and ${\rm ECN}(6_{\{3\}},4),$ the position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ can be considered as a position $(n_{0}+n_{3},n_{1}+n_{4},n_{2}+n_{5})$ in three-pile nim. Therefore, $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position if and only if $(n_{0}+n_{3})\oplus(n_{1}+n_{4})\oplus(n_{2}+n_{5})=0$.
@@
Let $N$ be the set of positions which are not in $P$. It is clear that the terminal position is in $P$.
@@
Lastly, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in N$. Let $k=\min(n_{0}\oplus n_{3},n_{1}\oplus n_{4},n_{2}\oplus n_{5})$. Without loss of generality, assume that $k=n_{0}\oplus n_{3}$. Then, since $M\not\in P$, at least one of $n_{1}\oplus n_{4}>k$ and $n_{2}\oplus n_{5}>k$ holds. In addition, each of $(n_{1}),(n_{4}),(n_{2}),(n_{5}),(n_{1},n_{2}),(n_{1},n_{5}),(n_{4},n_{2})$ and $(n_{4},n_{5})$ can be reduced in one move in ${\rm ECN}(6_{\{1,2\}},2)$. Therefore, $M$ has an option $M^{\prime}$ that satisfies $M^{\prime}=(n_{0},n^{\prime}_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n^{\prime}_{5})$, where $n_{0}\oplus n_{3}=n^{\prime}_{1}\oplus n^{\prime}_{4}=n^{\prime}_{2}\oplus n^{\prime}_{5}$. ∎
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}\}$.
@@
Consider a position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. This position is a selective sum of $(n_{0},n_{3})$, $(n_{1},n_{4})$, and $(n_{2},n_{5})$, which are positions of two-pile nim. Therefore, $n_{0}\oplus n_{3}=n_{1}\oplus n_{4}=n_{2}\oplus n_{5}=0$, that is, $n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}$ is a necessary and sufficient condition for that $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position.
@@
Let $P$ be the set of positions in ${\rm ECN}(6_{\{1,2\}},4)$ that satisfies the following:
  • Assume that $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. When each $n_{i}$ is expanded in the binary form, the summand of each digit contains four or no $1$s. In addition, for each digit contains four $1$s, the $1$s are in $(n_{i},n_{(i+1)\bmod 6},n_{(i+3)\bmod 6},\allowbreak n_{(i+4)\bmod 6})$ or $(n_{i},n_{(i+1)\bmod 6},n_{(i+2)\bmod 6},n_{(i+4)\bmod 6})$ for an integer $i\in\{0,1,2,3,4,5\}$.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},4)$. Then, M is a $\mathcal{P}$-position if and only if $M\in P$.
@@
We prove this theorem by using Theorem 3. Note that $\{\{v_{i},v_{(i+1)\bmod 6},v_{(i+3)\bmod 6},v_{(i+4)\bmod 6}\}\mid 0\leq i\leq 5\}=\{\{v_{i},v_{(i+1)},v_{(i+3)},v_{(i+4)\bmod 6}\}\mid 0\leq i\leq 2\}.$ Let $\mathcal{B}=\{\emptyset\}\cup\{\{v_{i},v_{(i+1)\bmod 6},v_{(i+2)\bmod 6},v_{(i+4)\bmod 6}\}\mid 0\leq i\leq 5\}\cup\{\{v_{i},v_{(i+1)},v_{(i+3)},v_{(i+4)\bmod 6}\}\mid 0\leq i\leq 2\}$. We show $\mathcal{B}$ satisfies the three conditions of nim-basis.
For (A), it is clear that $\emptyset\in\mathcal{B}$.
For (B), assume that $F\in\Delta$ and $B,B^{\prime}\in\mathcal{B}$ satisfy $B^{\prime}=F\sqcup B$. When $B=\emptyset,$ since $B^{\prime}=F$ and $F\not\in\mathcal{B}$ if $|F|>0$, $B^{\prime}=F=\emptyset$ holds. For the other cases, since each of $B$ and $B^{\prime}$ has four vertices, $B=B^{\prime}$ and $F$ must be the empty face.
For (C), assume that $F\in\Delta$ and $S$ is a subset of $V$. We show that there exist faces $K,G\in\Delta$, with $K\subseteq F\subseteq G,$ such that $G-F\subseteq S$ and $(S-G)\sqcup K\in\mathcal{B}$.
If $S\cup F$ is a face of $\Delta$, then let $G=S\cup F,K=\emptyset$. We have $(G-F)\subseteq S$ and $(S-G)\sqcup K=\emptyset\in\mathcal{B}$.
If $S\cup F$ is not a face of $\Delta$, we consider three cases. Note that if a subset of $V$ has less than four vertices, it is a face of $\Delta.$ If $S\cup F$ is a face of $\Delta,$ we consider the following three cases.
  • (i)
    When $S\cup F$ has four vertices and is not a face of $\Delta$, we note that $S\cup F\in\mathcal{B}$. Let $G=K=F$. Then, $G-F=\emptyset\subseteq S$ and $(S-G)\sqcup K=S\cup F\in\mathcal{B}$.
  • (ii)
    When $S\cup F$ has five vertices, without loss of generality, we assume that $S\cup F=\{v_{0},v_{1},v_{2},v_{3},v_{4}\}$. We consider five cases:
    • (ii-a)
      When $F=\{v_{0},v_{4}\}$, let $G=F\cup\{v_{2}\}$ and $K=F$. Then, $G-F=\{v_{2}\}\subset S$ and $(S-G)\sqcup K=\{v_{0},v_{1},v_{3},v_{4}\}\in\mathcal{B}$.
    • (ii-b)
      When $F(\neq\{v_{0},v_{4}\})$ has two, three, or four vertices, let $G=F$. Then, we have $G-F=\emptyset\subseteq S$. Since $F$ has at least two vertices, SG=SFS-G=S-F has at most three vertices.
      If $v_{1}\in F,$ SGS-G is a subset of $\{v_{0},v_{2},v_{3},v_{4}\}$, then $K=F-\{v_{1}\}$ satisfies $K\subset F$ and $(S-G)\sqcup K=S\cup F-\{v_{1}\}=\{v_{0},v_{2},v_{3},v_{4}\}\in\mathcal{B}$.
      Similarly, if $v_{2}\in F$ (resp. $v_{3}\in F),$ then $K=F-\{v_{2}\}$ (resp. $K=F-\{v_{3}\}$) satisfies $K\subset F$ and $(S-G)\sqcup K\in\mathcal{B}$.
      Since $F$ has at least two vertices, the remaining case is $\{v_{0},v_{4}\}\subset F$. From the rule of ${\rm ECN}(6_{\{1,2\}},4)$ and $v_{5}\not\in F$, $F=\{v_{0},v_{2},v_{4}\}$ or $F=\{v_{0},v_{4}\}$ holds, but the latter case has been considered. For the former case, let $K=F-\{v_{2}\}$. We have $K\subset F$ and $(S-G)\sqcup K=S\cup F-\{v_{2}\}=\{v_{0},v_{1},v_{3},v_{4}\}\in\mathcal{B}$.
    • (ii-c)
      When $F=\{v_{1}\},F=\{v_{2}\},$ or $F=\{v_{3}\}$, let $G=F$ and $K=\emptyset$. Then, $G-F=\emptyset\subseteq S$ and $(S-G)\sqcup K=S-G\in\{\{v_{0},v_{2},v_{3},v_{4}\},\{v_{0},v_{1},v_{3},v_{4}\},\{v_{0},v_{1},v_{2},v_{4}\}\}\subset\mathcal{B}$.
    • (ii-d)
      When $F=\{v_{0}\}$ (resp. $F=\{v_{4}\}$), let $G=\{v_{0},v_{1}\}$ (resp. $G=\{v_{3},v_{4}\}$) and $K=F$. Then, $G-F=\{v_{1}\}\subseteq S$ (resp. $G-F=\{v_{3}\}\subseteq S$) and $(S-G)\sqcup K=\{v_{0},v_{2},v_{3},v_{4}\}\in\mathcal{B}$ (resp. $(S-G)\sqcup K=\{v_{0},v_{1},v_{2},v_{4}\}\in\mathcal{B}$).
    • (ii-e)
      When $F=\emptyset$, let $G=\{v_{2}\}$ and $K=F=\emptyset$. Then, $G-F=\{v_{2}\}\subseteq S$ and $(S-G)\sqcup K=\{v_{0},v_{1},v_{3},v_{4}\}\in\mathcal{B}$.
  • (iii)
    When $S\cup F$ has six vertices, without loss of generality, we assume that $F\in\{\emptyset,\{v_{0}\},\{v_{0},v_{1}\},\{v_{0},v_{2}\},\allowbreak\{v_{0},v_{3}\},\{v_{0},v_{1}.v_{2}\},\{v_{0},v_{1},v_{3}\},\{v_{0},v_{2},v_{4}\},\{v_{0},v_{1},v_{2},v_{3}\}\}.$ We consider four cases:
    • (iii-a)
      When $F=\emptyset,F=\{v_{0}\}$, or $F=\{v_{0},v_{2}\}$, let $G=\{v_{0},v_{2}\}$ and $K=\emptyset$. Then, $F\subseteq G$ and $(G-F)\subseteq S$. Also, $K\subseteq F$ and $(S-G)\sqcup K=\{v_{1},v_{3},v_{4},v_{5}\}\in\mathcal{B}$ hold.
    • (iii-b)
      When $F=\{v_{0},v_{1}\},$ let $G=\{v_{0},v_{1},v_{2}\}$ and $K=\{v_{1}\}$. Then, $F\subseteq G$ and $(G-F)=\{v_{2}\}\subset S$. Also, $K\subseteq F$ and $(S-G)\sqcup K=\{v_{1},v_{3},v_{4},v_{5}\}\in\mathcal{B}$ hold.
    • (iii-c)
      When $F=\{v_{0},v_{3}\},F=\{v_{0},v_{1},v_{3}\},$ or $F=\{v_{0},v_{1},v_{2},v_{3}\}$, let $G=F$ and $K=\{v_{1},v_{2},v_{4},v_{5}\}\cap F\subseteq F$. Note that SG=SFS-G=S-F is a subset of $\{v_{1},v_{2},v_{4},v_{5}\}\in\mathcal{B}$. Then, $G-F=\emptyset\subset S$ and $(S-G)\sqcup K=(S-F)\sqcup(\{v_{1},v_{2},v_{4},v_{5}\}\cap F)=\{v_{1},v_{2},v_{4},v_{5}\}\in\mathcal{B}$ hold.
    • (iii-d)
      When $F=\{v_{0},v_{1},v_{2}\}$ or $F=\{v_{0},v_{2},v_{4}\},$ let $G=F$ and $K=\{v_{1},v_{3},v_{4},v_{5}\}\cap F\subseteq F$. Note that SG=SFS-G=S-F is a subset of $\{v_{1},v_{3},v_{4},v_{5}\}.$ Then, $G-F=\emptyset\in S$ and $(S-G)\sqcup K=(S-F)\sqcup(\{v_{1},v_{3},v_{4},v_{5}\}\cap F)=\{v_{1},v_{3},v_{4},v_{5}\}\in\mathcal{B}.$
For ${\rm ECN}(6_{\{1,3\}},2),$ we show the following more generalized theorem.
@@
Proof.
In ${\rm ECN}((2m)_{\{1,3,\ldots,h\}},2)$, the player can reduce at most one of $n_{0},n_{2},\ldots,n_{2m-2}$ and at most one of $n_{1},n_{3},\ldots,n_{2m-1}$ in one move. Therefore, this ruleset is the selective sum of two positions $(n_{0},n_{2},\ldots,n_{2m-2})$ and $(n_{1},n_{3},\ldots,n_{2m-1})$ in $m$-pile nim. Thus, from Theorems 1 and 6, $M$ is a $\mathcal{P}$-position if and only if $n_{0}\oplus n_{2}\oplus\cdots\oplus n_{2m-2}=n_{1}\oplus n_{3}\oplus\cdots\oplus n_{2m-1}=0.$
Corollary 1.
@@
Note that Theorem 11 is also a generalization of the result on ${\rm ECN}(4_{\{1\}},2)$.
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{2,3\}},3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{2}+n_{4}=n_{1}+n_{3}+n_{5},n_{0}\oplus n_{1}\oplus n_{2}=0,n_{0}\leq n_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}\}$.
@@
Next, consider a position $M\in_{\circlearrowleft}P$. Without loss of generality, we assume that $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. Then, $n_{0}\leq n_{3},n_{1}\leq n_{4},$ and $n_{2}\leq n_{5}$ hold. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5})$ be an option of $M$. When the player reduces some of $n_{0},n_{2},n_{4}$, or reduces some of $n_{1},n_{3},n_{5}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}\neq n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ holds and $M^{\prime}\not\in_{\circlearrowleft}P$. For the case that the player reduces the same number from $n_{0}$ and $n_{3}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ is satisfied but $n^{\prime}_{0}\oplus n^{\prime}_{1}\oplus n^{\prime}_{2}=n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$ holds. We show $M^{\prime}\not\in_{\circlearrowleft}P$ by contradiction:
@@
    Assume that $n^{\prime}_{2}\oplus n^{\prime}_{3}\oplus n^{\prime}_{4}=n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=0$. In order to satisfy $M^{\prime}\in_{\circlearrowleft}P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4}$ hold. However, then we have $n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=n_{2}\oplus n^{\prime}_{0}\oplus n_{1}=0$, which contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.
@@
    Assume that $n^{\prime}_{3}\oplus n^{\prime}_{4}\oplus n^{\prime}_{5}=n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0$. In order to satisfy $M^{\prime}\in_{\circlearrowleft}P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1},n_{5}\leq n_{2}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4},n_{2}=n_{5}$ hold. This contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.